OpenStudy (anonymous):

The circumference of a circle is 50.24 centimeters. Find the area of the circle. (Use 3.14 for π.)

OpenStudy (anonymous):

50.24 cm2 803. 84 cm2 100.48 cm2 200.96 cm it would be one of these ansers

OpenStudy (campbell_st):

you need to find the radius of the circle. Use the curcumference of circle formula \[C = 2\pi r\] in your question you have \[50.24 = 3 \times 3.14 \times r\] you need to solve for r Once you have found the radius then use the area of a circle formula \[A=\pi r^2\]

OpenStudy (anonymous):

im lost lol

OpenStudy (campbell_st):

ok... can you find the value of the radius from the circumference information 50.24 = 2 x 3.14 x r

OpenStudy (anonymous):

yea go on

OpenStudy (campbell_st):

what did you get for r...?

OpenStudy (anonymous):

idk thats what i ask for you for lol \

OpenStudy (anonymous):

im not good at math at all

OpenStudy (campbell_st):

well 1st divide 50.24 by 2 what do you get..?